Situation 2: CVP Analysis (15 Points) Aranguez Corporation produces a molded plastic casing, LX20A, for desktop computers. Summary data from its 2013 income statement are as follows: Variable costs ixed costs (3,600,000 Jane Dall, Aranguezs president, is very concerned about Aranguez Corporations poor profitability. She asks Giselle James, production manager, and Lester Saine, controller, to see if there are ways to reduce costs. After two weeks, James returns with a proposal to reduce variable costs to 50% of revenues by reducing the costs Aranguez currently incurs for safe disposal of waste plastic. Saline is concerned that this would expose the company to potential environmental liabilities. He tells James, We would need to estimate some of these potential environmental costs and include them in our analysis. You cannot do that, James replies. We are not violating any laws. There is some possibility that we may have to incur environmental costs in the future, but if we bring it up now, this proposal will not go through because our senior management always assumes these costs to be larger than they turn out to be. The market is very tough, and we are in danger of shutting down the c colleagues to lose their jobs. The only reason our competitors are making money is because they are doing exactly what I am p ompany. We dont want all our roposing. Required . Calculate Aranguez Corporations breakeven revenues for 2013. (2 Points) 2. Calculate Aranguez Corporations breakeven revenues for 2013, if variable costs were 50% of revenues. (2 Points) 3. Calculate Aranguez Corporations margin of safety as a percent of sales assuming that variable costs have been reduced to 50% of sales. (2 Points) 4. Given Giselle Jamess comments, what should Lester Saline do? Be sure to refer to the IMAs Ethical Guidance which is presented on page 864 of the course text. 9 Points)
Answer #1
Requirement 1
Breakeven revenues for 2013
Revenues        6,000,000
Variable cost      (3,600,000)
Contribution margin        2,400,000
Fixed costs (2,000,000)
Operating income            400,000
Contribution margin ratio = Contribution margin /Sales X100
                 =2,400,000/6,000,000 X100
                 =40%
Breakeven revenues = Fixed costs/ CM ratio
               = 2,000,000 /40%
               = 5,000,000
Requirement 2
Breakeven revenues for 2013, if variable costs are 50% of revenues
Revenues        6,000,000
Variable cost      (3,000,000)
Contribution margin        3,000,000
Fixed costs (2,000,000)
Operating income        1,000,000
Contribution margin ratio = Contribution margin /Sales X100
                 =3,000,000/6,000,000 X100
                 =50%
Breakeven revenues = Fixed costs/ CM ratio
               = 2,000,000 /50%
               = 4,000,000
Requirement 3
Margin of safety as a percentage of sales
          = (Actual sales - break even sales)/ Actual sales X100
          = (6,000,000 -4,000,000)/ 6,000,000 X100
          = 33.33%

Requirement 4

I dont have data to refer the IMA 's Ethical guidance